# On the singular loci of higher secant varieties of Veronese embeddings

### v2

@@ 4.2. Singularity
By the conditions in the Part (iii) of these two Theorems, we see that $k-1\geq\lceil\frac{{m+d\choose m}}{m+1}\rceil$ if $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ is never defective or $k-1\geq\lceil\frac{{m+d\choose m}}{m+1}\rceil+1$ else if $(d,m)$ appears in the defective list of Alexander-Hiroschowitz. In any case, we have $\sigma_{k-1}(v_{d}({\mathbb{P}}^{m}))=\langle\,v_{d}({\mathbb{P}}^{m})\,\rangle$. Hence, $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))=\sigma_{k-1}(v_{d}({\mathbb{P}}^{m}))\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ and the assertion follows. ∎
@@
It is shown in [15] by giving the defining equations of $\sigma_{4}(v_{3}({\mathbb{P}}^{3}))$ explicitly. ∎
@@ 5. Case of 4-th secant variety of Veronese embedding
In this section we aim to prove Theorem 5 as an investigation of the singular loci of the fourth secant variety (i.e. $k=4$) of any Veronese variety. This theorem consists of one part dealing with (non-)singularity of points in full-secant loci (i.e. $m=3$) and the other part for points in the maximum subsecant loci $\Sigma_{4,d}(\min\{k-1,n\}-1)$. So, we will obtain Theorem 5 by proving Theorem 29 (Part (i) of Theorem 5) and Corollary 30 (Part (ii) and (iii) of Theorem 5).
@@ 5.1. Equations by Young flattening
In [23] another source of equations for secant varieties of Veronese was introduced via so-called Young flattening. Here we briefly review construction of a certain type of Young flattening and exploit it for computation of conormal space of a given form.
@@
Thus, from $k\binom{n}{a}+1$ minors of the matrix $\mathrm{YF}^{a}_{d_{1},d_{2},n}(f)$ we obtain a set of equations for $\sigma_{k}(v_{d}({\mathbb{P}}V))$ for this range of $k$ (for some values of $k,d,d^{\prime},a$, it is known that these minors cut $\sigma_{k}(v_{d}({\mathbb{P}}V))$ as an irreducible component (see [23, theorem 1.2.3])).
@@
$\hat{N}^{\ast}_{f}\sigma_{k}(v_{d}({\mathbb{P}}V))\supseteq(\ker\mathrm{YF}_{d_{1},d_{2},n}^{a}(f))\cdot(\operatorname{im}\mathrm{YF}_{d_{1},d_{2},n}^{a}(f))^{\perp}~{},$
where the right hand side is to be understood as the image of the multiplication $S^{d_{1}}V^{\ast}\otimes\bigwedge^{a}V\otimes S^{d_{2}}V^{\ast}\otimes\bigwedge^{a+1}V^{\ast}\to S^{d}V^{\ast}$.
@@
This proposition comes directly from the same idea as Proposition 23 by applying it to a linear embedding $S^{d}V\hookrightarrow S^{d_{1}}V\otimes\bigwedge^{a}V^{\ast}\otimes S^{d_{2}}V\otimes\bigwedge^{a+1}V$. Since $\operatorname{rank}\mathrm{YF}_{d_{1},d_{2},n}^{a}(f)=k{n\choose a}$ and, as observed before, $v_{d}({\mathbb{P}}V)$ is contained in $\sigma_{n\choose a}(Seg({\mathbb{P}}(S^{d_{1}}V\otimes\bigwedge^{a}V^{\ast})\times{\mathbb{P}}(S^{d_{2}}V\otimes\bigwedge^{a+1}V)))$ in ${\mathbb{P}}(S^{d_{1}}V\otimes\bigwedge^{a}V^{\ast}\otimes S^{d_{2}}V\otimes\bigwedge^{a+1}V)$ and not in the previous secants of the same Segre variety, it is straightforward from the proof of Proposition 23 (i.e. $p={n\choose a}$ case). ∎
@@ 5.2. Singularity and non-singularity
whose fibers $\pi^{-1}({\mathbb{P}}U)$’s are all isomorphic to $\sigma_{4}(v_{d}({\mathbb{P}}U))\setminus(\Sigma_{4,d}(2;{{\mathbb{P}}U})\cup\sigma_{3}(v_{d}({\mathbb{P}}U)))$, where recall that $\Sigma_{4,d}(2;{{\mathbb{P}}^{n}})\subset\sigma_{4}(v_{d}({\mathbb{P}}^{n}))$ is the maximum subsecant locus, i.e. the union of all $\sigma_{4}(v_{d}({\mathbb{P}}^{1}))$’s and $\sigma_{4}(v_{d}({\mathbb{P}}^{2}))$’s in $\sigma_{4}(v_{d}({\mathbb{P}}^{n}))$. So, we can reduce the proof of theorem to the case of $n=3$.
In case of $n=3$, there is a list of normal forms in $\sigma_{4}(v_{d}({\mathbb{P}}^{3}))\setminus(\Sigma_{4,d}(\min\{k-1,n\}-1)\cup\sigma_{3}(v_{d}({\mathbb{P}}^{3})))$ due to Landsberg-Teitler (see [21, theorem 10.9.3.1]) such as
$f_{1}={x_{0}}^{d}+{x_{1}}^{d}+{x_{2}}^{d}+{x_{3}}^{d}$
$f_{2}={x_{0}}^{d-1}x_{1}+{x_{2}}^{d}+{x_{3}}^{d}$
$f_{3}={x_{0}}^{d-1}x_{1}+{x_{2}}^{d-1}{x_{3}}$
$f_{4}={x_{0}}^{d-2}{x_{1}}^{2}+{x_{0}}^{d-1}x_{2}+{x_{3}}^{d}$
@@
Case (ii) $f_{2}={x_{0}}^{d-1}x_{1}+{x_{2}}^{d}+{x_{3}}^{d}$. Say $U={\mathbb{C}}\langle x_{0},x_{1},x_{2},x_{3}\rangle$. Consider the Young flattening $\mathrm{YF}_{d-2,1,3}^{1}(f_{2})\in S^{d-2}U\otimes U\otimes U^{\ast}\otimes\wedge^{2}U\simeq\operatorname{Hom}(S^{d-2}U^{\ast}\otimes U,U\otimes\wedge^{2}U)$ defined in (24). For simplicity, call this type of Young flattening by $\phi$ throughout the proof. Then, $\phi(f_{2})$ is
@@
for some nonzero $\alpha,\beta,\gamma,\delta,\epsilon\in{\mathbb{C}}$. Note that as a linear map $S^{d-2}U^{\ast}\otimes U\to U\otimes\wedge^{2}U$, $\operatorname{rank}\phi(x_{0}^{5})=3$ and $\operatorname{rank}\phi(f_{2})=4\cdot 3=12$. Thus, by Proposition 28 $(\ker\phi(f_{2}))\cdot(\operatorname{im}\phi(f_{2}))^{\perp}$ produces a subspace of $\hat{N}^{\ast}_{f_{2}}\sigma_{4}(v_{d}({\mathbb{P}}^{3}))$.
For $d=3$, the expected dimension of $\hat{N}^{\ast}_{f_{2}}\sigma_{4}(v_{3}({\mathbb{P}}^{3}))$ for the smoothness is (${3+3\choose 3}-16=4$ and the corresponding four can be chosen as $y_{0}y_{2}y_{3},y_{1}^{2}y_{2},y_{1}y_{2}y_{3},y_{1}^{2}y_{3}$ in ${S}^{3}U^{\ast}$ which are given by the product of $\{y_{1}\otimes x_{0},y_{2}\otimes x_{2},y_{3}\otimes x_{3}\}$ in $\ker\phi(f_{2})\subset U^{\ast}\otimes U$ and
@@
For any $d\geq 4$, in $\ker\phi(f_{2})\subset S^{d-2}U^{\ast}\otimes U$, one can find a subspace generated by $\{F\otimes x_{i}~{}|~{}F\in J_{d-2},~{}i=0,\ldots,3\}$, where $J=\langle y_{0}y_{2},y_{0}y_{3},y_{1}^{2},y_{1}y_{2},y_{1}y_{3},y_{2}y_{3}\rangle$ is an ideal in ${S}^{\bullet}U^{\ast}$. Also, in $\operatorname{im}\phi(f_{2})^{\perp}\subset U^{\ast}\otimes\wedge^{2}U^{\ast}$ there exists the same subspace as in (25). In this case, our $(\ker\phi(f_{2}))\cdot(\operatorname{im}\phi(f_{2}))^{\perp}$ contains the subspace of ${S}^{d}U^{\ast}$ generated by
@@
$\cup\{y_{1}^{d},\ldots,y_{1}^{2}y_{2}^{d-2},\ldots,y_{1}^{2}y_{3}^{d-2}\}\cup\{y_{1}y_{2}^{d-2}y_{3},\ldots,y_{1}y_{2}y_{3}^{d-2}\}\cup\{y_{2}^{d-2}y_{3},\ldots,y_{2}^{2}y_{3}^{d-2}\}$
(note that the terms above are listed in the lexicographical order). These monomial generators can be also represented as
($\left(\{y_{0}^{d},y_{0}^{d-1}y_{1},\cdots,y_{0}y_{3}^{d-1}\}\setminus\{y_{0}^{d},y_{0}^{d-1}y_{1},y_{0}^{d-1}y_{2},y_{0}^{d-1}y_{3},y_{0}^{d-2}y_{1}^{2},y_{0}^{d-2}y_{1}y_{2},y_{0}^{d-2}y_{1}y_{3},y_{0}^{d-3}y_{1}^{3},y_{0}y_{2}^{d-1},y_{0}y_{3}^{d-1}\}\right)$
($\cup\left(\{y_{1}^{d},y_{1}^{d-1}y_{2},\cdots,y_{3}^{d}\}\setminus\{y_{1}y_{2}^{d-1},y_{1}y_{3}^{d-1},y_{2}^{d},y_{2}^{d-1}y_{3},y_{2}y_{3}^{d-1},y_{3}^{d}\}\right)~{},$
@@
$\dim\hat{N}^{\ast}_{f_{2}}\sigma_{4}(v_{d}({\mathbb{P}}^{3}))$ $\geq\dim(\ker\phi(f_{2}))\cdot(\operatorname{im}\phi(f_{2}))^{\perp}~{},$
{($\geq\bigg{\{}{d-1+3\choose 3}-10+{d+2\choose 2}-6\bigg{\}}=\binom{d+3}{3}-16~{}.$
@@
$~{}y_{2}\otimes y_{0}\wedge y_{1},~{}y_{2}\otimes y_{0}\wedge y_{1},~{}y_{3}\otimes y_{0}\wedge y_{1},~{}y_{3}\otimes y_{0}\wedge y_{3},~{}y_{3}\otimes y_{1}\wedge y_{3}~{}\rangle~{}$
@@
Similarly, for the case of $d\geq 4$, $(\ker\phi(f_{3}))\cdot(\operatorname{im}\phi(f_{3}))^{\perp}$ contains a subspace of $\hat{N}^{\ast}_{f_{3}}\sigma_{4}(v_{d}({\mathbb{P}}^{3}))\subset{S}^{d}U^{\ast}$ which is generated by
($\left(\{y_{0}^{d},y_{0}^{d-1}y_{1},\cdots,y_{0}y_{3}^{d-1}\}\setminus\{y_{0}^{d},y_{0}^{d-1}y_{1},y_{0}^{d-1}y_{2},y_{0}^{d-1}y_{3},y_{0}^{d-2}y_{1}^{2},y_{0}^{d-2}y_{1}y_{2},y_{0}^{d-2}y_{1}y_{3},y_{0}^{d-3}y_{1}^{3},y_{0}y_{2}^{d-1},y_{0}y_{2}^{d-2}y_{3}\}\right)$
($\cup\left(\{y_{1}^{d},y_{1}^{d-1}y_{2},\cdots,y_{3}^{d}\}\setminus\{y_{1}y_{2}^{d-1},y_{1}y_{2}^{d-2}y_{3},y_{2}^{d},y_{2}^{d-1}y_{3},y_{2}^{d-2}y_{3}^{2},y_{2}^{d-3}y_{3}^{3}\}\right)~{},$
using a subspace $\langle~{}\{F\otimes x_{i}~{}|~{}F\in J_{d-2},~{}i=0,\ldots,3\}~{}\rangle$ in $\ker\phi(f_{3})$, where $J$ is an ideal generated by $\{y_{0}y_{2},y_{0}y_{3},y_{1}^{2},y_{1}y_{2},y_{1}y_{3},y_{3}^{2}\}$ in ${S}^{\bullet}U^{\ast}$, and the same subspace in $\operatorname{im}\phi(f_{3})^{\perp}$ as (26). Thus, $\dim\hat{N}^{\ast}_{f_{3}}\sigma_{4}(v_{d}({\mathbb{P}}^{3}))\geq{d+3\choose 3}-16$, which means that $f_{3}$ is also smooth.
@@
and $\operatorname{rank}\phi(f_{4})=12$. $\hat{N}^{\ast}_{f_{4}}\sigma_{4}(v_{3}({\mathbb{P}}^{3}))$ contains a 4-dimensional subspace corresponding to $\langle-y_{0}y_{2}y_{3}+y_{1}^{2}y_{3},y_{1}y_{2}y_{3},y_{2}^{3},y_{2}^{2}y_{3}\rangle$ which can be spanned by $\{y_{2}\otimes x_{0},y_{3}\otimes x_{3}\}$ in $\ker\phi(f_{4})\subset U^{\ast}\otimes U$ and
@@
$+(d-1)(d-2)x_{0}^{d-3}x_{2}\otimes x_{0}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{0}\big{)}+d(d-1)x_{3}^{d-2}\otimes x_{3}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{3}\big{)}~{}.$
@@
$~{}-(d-1)y_{1}\otimes y_{1}\wedge y_{2}+2y_{2}\otimes y_{0}\wedge y_{2},~{}-2y_{0}\otimes y_{2}\wedge y_{3}+(d-1)y_{1}\otimes y_{1}\wedge y_{3}\}~{}.$
Then, one can check that $A_{1}\cdot B_{1}$ produces a subspace of $\hat{N}^{\ast}_{f_{4}}\sigma_{4}(v_{3}({\mathbb{P}}^{3}))$ in ${S}^{d}U^{\ast}$ which is the degree $d$-part of an ideal $I_{1}$ generated by 19 quartics
$\{~{}-4\underline{y_{0}^{2}y_{2}^{2}}+(2d+4)y_{0}y_{1}^{2}y_{2},~{}-2\underline{y_{0}^{2}y_{2}y_{3}}+3y_{0}y_{1}^{2}y_{3},~{}\underline{y_{0}^{2}y_{3}^{2}},~{}-2\underline{y_{0}y_{1}y_{2}^{2}}+(d-1)y_{1}^{3}y_{2},~{}-2\underline{y_{0}y_{1}y_{2}y_{3}}+3y_{1}^{3}y_{3},$
$~{}\underline{y_{0}y_{1}y_{3}^{2}},~{}-2\underline{y_{0}y_{2}^{3}}+(d-1)y_{1}^{2}y_{2}^{2},~{}-2\underline{y_{0}y_{2}^{2}y_{3}}+(d-1)y_{1}^{2}y_{2}y_{3},~{}-2\underline{y_{0}y_{2}y_{3}^{2}}+(d-1)y_{1}^{2}y_{3}^{2},~{}\underline{y_{1}^{3}y_{3}},$
$~{}\underline{y_{1}^{2}y_{2}^{2}},~{}\underline{y_{1}^{2}y_{2}y_{3}},~{}\underline{y_{1}^{2}y_{3}^{2}},~{}\underline{y_{1}y_{2}^{3}},~{}\underline{y_{1}y_{2}^{2}y_{3}},~{}\underline{y_{1}y_{2}y_{3}^{2}},~{}\underline{y_{2}^{4}},~{}\underline{y_{2}^{3}y_{3}},~{}\underline{y_{2}^{2}y_{3}^{2}}\}$
@@
=($={d+3\choose 3}-16\quad(d\geq 4)~{}.$
This implies that (${d+3\choose 3}-16\geq\dim\hat{N}^{\ast}_{f_{4}}\sigma_{4}(v_{d}({\mathbb{P}}^{3}))\geq H(I,d)={d+3\choose 3}-16$, which means that $\sigma_{4}$ is also smooth at $f_{4}$.

### v3

@@ 4.2. Singularity
By the conditions in the Part (iii) of these two theorems, we see that
$k-1\geq\Bigl{\lceil}\tfrac{{m+d\choose m}}{m+1}\Bigr{\rceil}$
if $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ is never defective, or
$k-1\geq\Bigl{\lceil}\tfrac{{m+d\choose m}}{m+1}\Bigr{\rceil}+1$
else if $(d,m)\in\{(3,4),(4,2),(4,3),(4,4)\}$, the defective list of Alexander-Hiroschowitz. In any case, we have $\sigma_{k-1}(v_{d}({\mathbb{P}}^{m}))=\langle\,v_{d}({\mathbb{P}}^{m})\,\rangle$. Hence,
$\sigma_{k}(v_{d}({\mathbb{P}}^{m}))=\sigma_{k-1}(v_{d}({\mathbb{P}}^{m}))\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$
and the assertion follows. ∎
@@
It is shown in [FH24] by giving the defining equations of $\sigma_{4}(v_{3}({\mathbb{P}}^{3}))$ explicitly. ∎
@@ 5. Case of 4-th secant variety of Veronese embedding
In this section we aim to prove Theorem 5 as an investigation of the singular loci of the fourth secant variety (i.e., $k=4$) of any Veronese variety. This theorem consists of one part dealing with the (non-)singularity of points in full-secant loci (i.e., $m=3$) and the other part for points in the maximum subsecant loci $\Sigma_{4,d}(\min\{k-1,n\}-1)$. So, we will obtain Theorem 5 by proving Theorem 29 (Part (i) of Theorem 5) and Corollary 30 (Part (ii) and (iii) of Theorem 5).
@@ 5.1. Equations by Young flattening
In [LO13] another source of equations for secant varieties of Veronese varieties was introduced via the so-called Young flattening. Here we briefly review the construction of a certain type of Young flattening and use it to compute the conormal space of a given form.
@@
Thus, from $k\binom{n}{a}+1$ minors of the matrix $\mathrm{YF}^{a}_{d_{1},d_{2},n}(f)$ we obtain a set of equations for $\sigma_{k}(v_{d}({\mathbb{P}}V))$ for this range of $k$ (for some values of $k,d,d^{\prime},a$, it is known that these minors cut $\sigma_{k}(v_{d}({\mathbb{P}}V))$ as an irreducible component (see [LO13, Theorem 1.2.3])).
@@
$\hat{N}^{\ast}_{[f]}\sigma_{k}(v_{d}({\mathbb{P}}V))\supseteq(\ker\mathrm{YF}_{d_{1},d_{2},n}^{a}(f))\cdot(\operatorname{im}\mathrm{YF}_{d_{1},d_{2},n}^{a}(f))^{\perp}~{},$
where the right hand side is to be understood as the image of the multiplication
$S^{d_{1}}V^{\ast}\otimes\bigwedge^{a}V\otimes S^{d_{2}}V^{\ast}\otimes\bigwedge^{a+1}V^{\ast}\to S^{d}V^{\ast}.$
@@
This proposition follows directly from the same idea as Proposition 23 by applying it to a linear embedding
$S^{d}V\hookrightarrow S^{d_{1}}V\otimes\bigwedge^{a}V^{\ast}\otimes S^{d_{2}}V\otimes\bigwedge^{a+1}V.$
Since $\operatorname{rank}\mathrm{YF}_{d_{1},d_{2},n}^{a}(f)=k{n\choose a}$ and, as observed before, $v_{d}({\mathbb{P}}V)$ is contained in
$\sigma_{n\choose a}(Seg({\mathbb{P}}(S^{d_{1}}V\otimes\bigwedge^{a}V^{\ast})\times{\mathbb{P}}(S^{d_{2}}V\otimes\bigwedge^{a+1}V)))\subset{\mathbb{P}}(S^{d_{1}}V\otimes\bigwedge^{a}V^{\ast}\otimes S^{d_{2}}V\otimes\bigwedge^{a+1}V)$
and not in the previous secants of the same Segre variety, it is straightforward from the proof of Proposition 23 (i.e., the case $p={n\choose a}$). ∎
@@ 5.2. Singularity and non-singularity
whose fibers $\pi^{-1}({\mathbb{P}}U)$’s are all isomorphic to $\sigma_{4}(v_{d}({\mathbb{P}}U))\setminus(\Sigma_{4,d}(2;{{\mathbb{P}}U})\cup\sigma_{3}(v_{d}({\mathbb{P}}U)))$, recalling that $\Sigma_{4,d}(2;{{\mathbb{P}}^{n}})\subset\sigma_{4}(v_{d}({\mathbb{P}}^{n}))$ is the maximum subsecant locus, i.e., the union of all $\sigma_{4}(v_{d}({\mathbb{P}}^{1}))$’s and $\sigma_{4}(v_{d}({\mathbb{P}}^{2}))$’s in $\sigma_{4}(v_{d}({\mathbb{P}}^{n}))$. So, we can reduce the proof of theorem to the case of $n=3$.
In case of $n=3$, there is a list of normal forms in $\sigma_{4}(v_{d}({\mathbb{P}}^{3}))\setminus(\Sigma_{4,d}(\min\{k-1,n\}-1)\cup\sigma_{3}(v_{d}({\mathbb{P}}^{3})))$ due to Landsberg-Teitler (see [Land12, Theorem 10.9.3.1]) such as
  1. (i)
    $f_{1}={x_{0}}^{d}+{x_{1}}^{d}+{x_{2}}^{d}+{x_{3}}^{d}$,
  2. (ii)
    $f_{2}={x_{0}}^{d-1}x_{1}+{x_{2}}^{d}+{x_{3}}^{d}$,
  3. (iii)
    $f_{3}={x_{0}}^{d-1}x_{1}+{x_{2}}^{d-1}{x_{3}}$ ,
  4. (iv)
    $f_{4}={x_{0}}^{d-2}{x_{1}}^{2}+{x_{0}}^{d-1}x_{2}+{x_{3}}^{d}$,
  5. (v)
@@
Case (ii) $f_{2}={x_{0}}^{d-1}x_{1}+{x_{2}}^{d}+{x_{3}}^{d}$. Say
$U={\mathbb{C}}\langle x_{0},x_{1},x_{2},x_{3}\rangle.$
Consider the Young flattening
$\mathrm{YF}_{d-2,1,3}^{1}(f_{2})\in S^{d-2}U\otimes U\otimes U^{\ast}\otimes\wedge^{2}U\simeq\operatorname{Hom}(S^{d-2}U^{\ast}\otimes U,U\otimes\wedge^{2}U)$
defined in (32). For simplicity, we will denote this type of Young flattening by $\phi$ throughout the proof. Then, $\phi(f_{2})$ is
@@
for some nonzero $\alpha,\beta,\gamma,\delta,\epsilon\in{\mathbb{C}}$. Note that as a linear map $S^{d-2}U^{\ast}\otimes U\to U\otimes\wedge^{2}U$, $\operatorname{rank}\phi(x_{0}^{5})=3$ and $\operatorname{rank}\phi(f_{2})=4\cdot 3=12$. Thus, by Proposition 28 $(\ker\phi(f_{2}))\cdot(\operatorname{im}\phi(f_{2}))^{\perp}$ produces a subspace of $\hat{N}^{\ast}_{[f_{2}]}\sigma_{4}(v_{d}({\mathbb{P}}^{3}))$.
For $d=3$, the expected dimension of $\hat{N}^{\ast}_{[f_{2}]}\sigma_{4}(v_{3}({\mathbb{P}}^{3}))$ for the smoothness is (${3+3\choose 3}-16=4$ and the corresponding four can be chosen as $y_{0}y_{2}y_{3},y_{1}^{2}y_{2},y_{1}y_{2}y_{3},y_{1}^{2}y_{3}$ in ${S}^{3}U^{\ast}$ which are given by the product of $\{y_{1}\otimes x_{0},y_{2}\otimes x_{2},y_{3}\otimes x_{3}\}$ in $\ker\phi(f_{2})\subset U^{\ast}\otimes U$ and
@@
For any $d\geq 4$, in $\ker\phi(f_{2})\subset S^{d-2}U^{\ast}\otimes U$, one can find a subspace generated by
$\{F\otimes x_{i}~{}|~{}F\in J_{d-2},~{}i=0,\ldots,3\},$
where $J=\langle y_{0}y_{2},y_{0}y_{3},y_{1}^{2},y_{1}y_{2},y_{1}y_{3},y_{2}y_{3}\rangle$ is an ideal in ${S}^{\bullet}U^{\ast}$. Also, in $\operatorname{im}\phi(f_{2})^{\perp}\subset U^{\ast}\otimes\wedge^{2}U^{\ast}$ there exists the same subspace as in (33). In this case, our $(\ker\phi(f_{2}))\cdot(\operatorname{im}\phi(f_{2}))^{\perp}$ contains the subspace of ${S}^{d}U^{\ast}$ generated by
$\{y_{0}^{2}y_{2}^{2},y_{0}^{2}y_{2}y_{3},y_{0}^{2}y_{3}^{2}\}\cup\{y_{0}y_{1}^{2}y_{2},\ldots,y_{0}y_{1}y_{3}^{2}\}\cup\{y_{0}y_{2}^{2}y_{3},y_{0}y_{2}y_{3}^{2}\}\cup\{y_{1}^{4},\ldots,y_{1}^{2}y_{3}^{2}\}\cup\{y_{1}y_{2}^{2}y_{3},y_{1}y_{2}y_{3}^{2},y_{2}^{2}y_{3}^{2}\}$
for $d=4$ and by
@@
$\cup\{y_{1}^{d},\ldots,y_{1}^{2}y_{2}^{d-2},\ldots,y_{1}^{2}y_{3}^{d-2}\}\cup\{y_{1}y_{2}^{d-2}y_{3},\ldots,y_{1}y_{2}y_{3}^{d-2}\}\cup\{y_{2}^{d-2}y_{3}^{2},\ldots,y_{2}^{2}y_{3}^{d-2}\}$
for any $d>4$ (note that the terms above are listed in the lexicographical order). In both cases, these monomial generators can be also represented as
($\bigl{(}\{y_{0}^{d},y_{0}^{d-1}y_{1},\cdots,y_{0}y_{3}^{d-1}\}\setminus\{y_{0}^{d},y_{0}^{d-1}y_{1},y_{0}^{d-1}y_{2},y_{0}^{d-1}y_{3},y_{0}^{d-2}y_{1}^{2},y_{0}^{d-2}y_{1}y_{2},y_{0}^{d-2}y_{1}y_{3},y_{0}^{d-3}y_{1}^{3},y_{0}y_{2}^{d-1},y_{0}y_{3}^{d-1}\}\bigr{)}$
($\cup\bigl{(}\{y_{1}^{d},y_{1}^{d-1}y_{2},\cdots,y_{3}^{d}\}\setminus\{y_{1}y_{2}^{d-1},y_{1}y_{3}^{d-1},y_{2}^{d},y_{2}^{d-1}y_{3},y_{2}y_{3}^{d-1},y_{3}^{d}\}\bigr{)}~{},$
@@
$\dim\hat{N}^{\ast}_{[f_{2}]}\sigma_{4}(v_{d}({\mathbb{P}}^{3}))$ $\geq\dim(\ker\phi(f_{2}))\cdot(\operatorname{im}\phi(f_{2}))^{\perp}$
{($\geq\bigg{\{}{d-1+3\choose 3}-10+{d+2\choose 2}-6\bigg{\}}=\binom{d+3}{3}-16.$
@@
$~{}y_{2}\otimes y_{0}\wedge y_{1},~{}y_{3}\otimes y_{0}\wedge y_{1},~{}y_{3}\otimes y_{0}\wedge y_{3},~{}y_{3}\otimes y_{1}\wedge y_{3}~{}\rangle~{}$
@@
Similarly, for the case of $d\geq 4$, $(\ker\phi(f_{3}))\cdot(\operatorname{im}\phi(f_{3}))^{\perp}$ contains a subspace of $\hat{N}^{\ast}_{[f_{3}]}\sigma_{4}(v_{d}({\mathbb{P}}^{3}))\subset{S}^{d}U^{\ast}$ which is generated by
($\bigl{(}\{y_{0}^{d},y_{0}^{d-1}y_{1},\cdots,y_{0}y_{3}^{d-1}\}\setminus\{y_{0}^{d},y_{0}^{d-1}y_{1},y_{0}^{d-1}y_{2},y_{0}^{d-1}y_{3},y_{0}^{d-2}y_{1}^{2},y_{0}^{d-2}y_{1}y_{2},y_{0}^{d-2}y_{1}y_{3},y_{0}^{d-3}y_{1}^{3},y_{0}y_{2}^{d-1},y_{0}y_{2}^{d-2}y_{3}\}\bigr{)}$
($\cup\bigl{(}\{y_{1}^{d},y_{1}^{d-1}y_{2},\cdots,y_{3}^{d}\}\setminus\{y_{1}y_{2}^{d-1},y_{1}y_{2}^{d-2}y_{3},y_{2}^{d},y_{2}^{d-1}y_{3},y_{2}^{d-2}y_{3}^{2},y_{2}^{d-3}y_{3}^{3}\}\bigr{)},$
using a subspace $\langle~{}\{F\otimes x_{i}~{}|~{}F\in J_{d-2},~{}i=0,\ldots,3\}~{}\rangle$ in $\ker\phi(f_{3})$, where $J$ is an ideal generated by $\{y_{0}y_{2},y_{0}y_{3},y_{1}^{2},y_{1}y_{2},y_{1}y_{3},y_{3}^{2}\}$ in ${S}^{\bullet}U^{\ast}$, and the same subspace in $\operatorname{im}\phi(f_{3})^{\perp}$ as (34). Thus, $\dim\hat{N}^{\ast}_{[f_{3}]}\sigma_{4}(v_{d}({\mathbb{P}}^{3}))\geq{d+3\choose 3}-16$, which means that $f_{3}$ is also smooth.
@@
and $\operatorname{rank}\phi(f_{4})=12$. $\hat{N}^{\ast}_{[f_{4}]}\sigma_{4}(v_{3}({\mathbb{P}}^{3}))$ contains a 4-dimensional subspace corresponding to $\langle-y_{0}y_{2}y_{3}+y_{1}^{2}y_{3},y_{1}y_{2}y_{3},y_{2}^{3},y_{2}^{2}y_{3}\rangle$ which can be spanned by $\{y_{2}\otimes x_{0},y_{3}\otimes x_{3}\}$ in $\ker\phi(f_{4})\subset U^{\ast}\otimes U$ and
@@
$+(d-1)(d-2)x_{0}^{d-3}x_{2}\otimes x_{0}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{0}\big{)}+d(d-1)x_{3}^{d-2}\otimes x_{3}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{3}\big{)}.$
@@
$~{}-(d-1)y_{1}\otimes y_{1}\wedge y_{2}+2y_{2}\otimes y_{0}\wedge y_{2},~{}-2y_{0}\otimes y_{2}\wedge y_{3}+(d-1)y_{1}\otimes y_{1}\wedge y_{3}\}.$
Then, one can check that $A_{1}\cdot B_{1}$ produces a subspace of $\hat{N}^{\ast}_{[f_{4}]}\sigma_{4}(v_{3}({\mathbb{P}}^{3}))$ in ${S}^{d}U^{\ast}$ which is the degree $d$-part of an ideal $I_{1}$ generated by 19 quartics
[…]
$\,-2\underline{y_{0}y_{1}y_{2}y_{3}}+3y_{1}^{3}y_{3},~{}\underline{y_{0}y_{1}y_{3}^{2}},~{}-2\underline{y_{0}y_{2}^{3}}+(d-1)y_{1}^{2}y_{2}^{2},~{}-2\underline{y_{0}y_{2}^{2}y_{3}}+(d-1)y_{1}^{2}y_{2}y_{3},~{}-2\underline{y_{0}y_{2}y_{3}^{2}}+(d-1)y_{1}^{2}y_{3}^{2},$
$\,~{}\underline{y_{1}^{3}y_{3}},~{}\underline{y_{1}^{2}y_{2}^{2}},~{}\underline{y_{1}^{2}y_{2}y_{3}},~{}\underline{y_{1}^{2}y_{3}^{2}},~{}\underline{y_{1}y_{2}^{3}},~{}\underline{y_{1}y_{2}^{2}y_{3}},~{}\underline{y_{1}y_{2}y_{3}^{2}},~{}\underline{y_{2}^{4}},~{}\underline{y_{2}^{3}y_{3}},~{}\underline{y_{2}^{2}y_{3}^{2}}\ \}$
@@
=($={d+3\choose 3}-16\quad(d\geq 4).$
This implies that
(${d+3\choose 3}-16\geq\dim\hat{N}^{\ast}_{[f_{4}]}\sigma_{4}(v_{d}({\mathbb{P}}^{3}))\geq H(I,d)={d+3\choose 3}-16,$
which means that $\sigma_{4}$ is also smooth at $f_{4}$.